# Rethinking Negative Instances for Generative Named Entity Recognition

### v1

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities for generalizing in unseen tasks. In the Named Entity Recognition (NER) task, recent advancements have seen the remarkable improvement of LLMs in a broad range of entity domains via instruction tuning, by adopting entity-centric schema. In this work, we explore the potential enhancement of the existing methods by incorporating negative instances into training. Our experiments reveal that negative instances contribute to remarkable improvements by (1) introducing contextual information, and (2) clearly delineating label boundaries. Furthermore, we introduce a novel and efficient algorithm named Hierarchical Matching, which is tailored to transform unstructured predictions into structured entities. By integrating these components, we present GNER, a Generative NER system that shows improved zero-shot performance across unseen entity domains. Our comprehensive evaluation illustrates our system’s superiority, surpassing state-of-the-art (SoTA) methods by 11 $F_{1}$ score in zero-shot evaluation. [^1] Code, datasets and models will be released at https://github.com/yyDing1/GNER.
@@ 1 Introduction
To calibrate the potential enhancement of including negative instances in training, we first conduct a preliminary study. We choose Flan-T5-large Chung et al. (2022) as the generative backbone model and design strategies for training with negative instances. Through experiments, we show that negative instances can significantly boost the model’s performance by (1) incorporating the contextual information, and (2) enhancing the label boundary between entities and non-entities. The possible drawback of introducing the negative instances is the increase of the prediction length, leading to inaccurate predictions, reflected by the word addition, omission and substitution. To tackle the inaccuracy drawbacks, we aim to design a more accurate and efficient algorithm to convert unstructured text into structured entities.
Inspired by the above observations, we design an effective and efficient Generative NER framework named GNER. We first design a proper task schema integrating negative instances into the instruction tuning process. Additionally, we design a Hierarchical Matching algorithm to tackle the issues in the structuring process efficiently. This innovation ensures accurate categorization and alignment of extracted entities. We also demonstrate that zero-shot performance can be enhanced with beam search through a self-correction mechanism. These strategic developments collectively advance the GNER framework, setting a new standard for accuracy and efficiency in the field of NER.
We conduct experiments on two representative generative models, Flan-T5 and LLaMA. The resulting models, GNER-T5 and GNER-LLaMA, outperform SoTA by a large margin. As stated in Fig. 1, GNER-LLaMA-7B outperforms the GoLLIE Sainz et al. (2023) trained on Code-LLaMA-7B by 8 $F_{1}$ score. Furthermore, compared to the similarly configured model UniversalNER, GNER-LLaMA-7B shows an improvement of 12.7 $F_{1}$ score, with a 2.5$\times$ boost in inference speed. Our GNER-T5-11B model also achieves SoTA performance in both zero-shot and supervised settings.
@@ 2 Related Work
Early works format Named Entity Recognition (NER) as a sequence labeling problem Chiu and Nichols (2016); Huang et al. (2015); Akbik et al. (2018); Qin et al. (2019). Among these, BERT Tagging Devlin et al. (2018) is the most representative one. Then, different methods are proposed to address more complex scenarios, i.e., nested and discontinuous NER. These methods regard NER as question answering Li et al. (2020a); Mengge et al. (2020), span classification Fu et al. (2021); Li et al. (2020b), dependency parsing Yu et al. (2020), word-level relation classification Li et al. (2022a), and so on. In most of these approaches, negative instances have played a crucial role in the training process, either by integrating all negative instances or employing sampling methods to select part of them Li et al. (2022b). However, the performance of the above-mentioned supervised models significantly decreases in zero-shot settings Liu et al. (2021), especially when the data and domain distribution significantly diverge from those seen of training.
@@
3 Preliminary Study
To explore possible improvements of recent entity-centric schema, we launch a preliminary exploration from two perspectives: (1) incorporating the context of entities into the training process and (2) enhancing the label boundary between entities and non-entities. We then discuss the accompanying longer sequence prediction issues to LLMs when presenting context and label boundaries.
3.1 Definitions
In this part, we give the formal definitions of the entity-centric method and negative instances. The Named Entity Recognition (NER) Task can be formally defined as a function mapping of input tokens $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and a pre-defined set of entity types $L=\{l_{1},l_{2},\ldots l_{m}\}$ to entity labels $Y=\{y_{1},y_{2},\ldots,y_{n}\}$. The positive and negative instances can be formulated as follows:
@@ 3.1 Definitions
where O represents non-entity text. Entity-centric generative methods train the model using $P$ as training instances. It is worth noting that the training instances mentioned here refer to tokens that directly influence the model’s weights through the backpropagation algorithm during training.
An example in the Preliminary Study Definition (section 3.1) Token inputs ($X$): John explored Tokyo , sampling its famed sushi , and flew back to New York . Entity type ($L$): [Person, Location]   Training prompt (section 3.2 and 3.3) w/o context: [John](Person) [Tokyo](Location) [New York](Location). w/ context length 2: [John](Person) explored [Tokyo](Location) , sampling …… back to [New York](Location) . w/ full context: [John](Person) explored [Tokyo](Location) , sampling its famed sushi , and flew back to [New York](Location). w/ full context and label boundary: John(B-Person) explored(O) Tokyo(B-Location) ,(O) sampling(O) its(O) famed(O) sushi(O) ,(O) and(O) flew(O) back(O) to(O) New(B-Location) York(I-Location) .(O)   Perdiction Issues (section 3.4) Omission and Addition: John(B-Person) first(O) explored(O) Tokyo(B-Location) ……(omitted text) and(O) then(O) flew(O) back(O) to(O) New(B-Location) York(I-Location) .(O) Substitution: John(B-Person) explored(O) Tokyo(B-Location) ,(O) sampling(O) its(O) famed(O) sushi(O) ,(O) and(O) returned(O) to(O) New(B-Location) York(I-Location) .(O)
Figure 3: Prompts that are used in our preliminary study.
@@ 3.2 Learning with Entity Context
The context surrounding entities plays a significant role in determining their categories. For example, phrases like “go to” and “travel to” are often followed by a Location entity. We integrate the contextual information before and after an entity into our training process to explicitly enable the model to recognize entities based on their surrounding context. To achieve this, we introduce negative instances that are closest to the entity, extending up to a length $L$, until encountering the boundary of the sentence, as part of our training instances. The remaining negative instances that are not selected as contextual information are represented by ellipses. An example of our constructed prompt used for training is shown in Fig. 3. From the example, we can also glimpse the importance of context, such as “flew to” guiding the model to pay more attention to the subsequent entity “New York”, rather than just remembering “New York” as a location in the training process. We conduct experiments to explore the impact of contextual length. For the experimental setup, we choose Flan-T5-large, a model with 783M parameters, as our backbone. Additionally, we sample 10K samples from the Pile-NER Zhou et al. (2023) dataset as our training set and 200 samples for each subtask of CrossNER Liu et al. (2021) validation set to evaluate the model’s zero-shot performance. Results are shown in Fig. 4. The improvement from a contextual length from 0 to 1 is particularly notable, suggesting that the closest negative instance to the entity plays a significant role in identifying the current entity. As the contextual length increases, performance progressively improves, showing that the model benefits from the context. Yet, beyond a certain point, the speed of improvement slows down, indicating the amount of contextual information available in the entity context is approaching saturation.
@@ 3.3 Entity Boundary of Generative Model
Method Strict $\text{Strict }F_{1}$ Partial $\text{Partial }F_{1}$ $\delta$
UniNER-7B 53.4 56.7 3.3
UniNER-7B w/ sup 61.8 64.8 3.0
Table 1: Zero-shot performance of UniversalNER Zhou et al. (2023). $\delta$ denotes the difference between strict match $F_{1}$ and partial match $F_{1}$, indicating the model’s ability to distinguish entity boundaries.
The above analysis has demonstrated the effectiveness of entity contexts. We then introduce label information of contexts to enhance the label boundary between entities and non-entities. To some extent, the results of Zhou et al. (2023) expose the problem of vague entity boundaries in model predictions, as shown by the significantly higher $F_{1}$ score for partial match compared to strict match. Specifically, strict match requires the boundary to exactly match the ground truth, while partial match considers predicted entities that overlap with but do not strictly match the ground truth as half correct, counting them as 0.5 in true positive. We list their results in Table 1. Motivated by this, we further enhance the label boundary between entities and non-entities by the following strategies: (1) for the non-entity part, we label the context surrounding an entity with the negative label “O”, (2) for the entity part, the beginning of an entity is marked as “B-”, and the inside of the entity is marked as “I-”. The training prompt and results are shown in Fig. 3 and Fig. 4, respectively. Compared to training with contexts alone, incorporating our enhanced label boundary strategies results in consistent improvements across various contextual lengths.
3.4 Long Sequence Prediction Issues
The prediction length of the model increases with the integration of entity contexts and label boundaries. A longer generation sequence might bring challenges to the popular generative LLMs, as illustrated by the example in Fig. 3. The model’s output may include omissions, additions, and substitutions of words. For instance, it may add conjunctions “first” and “then” to sentences, omit irrelevant non-entity text, such as “sampling its framed sushi”, or replace “flew back” with “returned”. We launch a detailed case study and find the potential causes of these issues: (1) noise in the original text, (2) missing words in the vocabulary, and (3) accumulative exposure bias. The representative examples, issue proportion, along with detailed analysis, are documented in Appendix C.
@@ 4 Method
In this section, we present our GNER framework. We start by describing our task schema, which integrates negative instances into the training process for better usage of contextual information (section 3.2) and sensitivity to the entity boundaries (section 3.3), followed by the correlated tuning strategies. In response to the issues in section 3.4, we propose an effective algorithm, named Hierarchical Matching, to convert the model’s unstructured text outputs into structured data, thereby enhancing the accuracy of our system.
@@ 4.1 Task Schema
ChatGPT - 52.4 39.8 66.6 68.5 67.0 5.3 32.8 47.5
${}^{\dagger}$ flan-t5-xxl (11B) 49.0 47.2 53.2 48.2 49.3 63.0 21.0 47.3
${}^{\dagger}$ Code-LLaMA-7B 59.1 62.7 67.8 57.2 55.5 63.0 43.4 58.4
${}^{\dagger}$ Code-LLaMA-13B 56.7 59.7 65.5 54.4 56.2 62.5 49.8 57.8
${}^{\ddagger}$ LLaMA-7B 53.5 59.4 65.0 60.8 61.1 42.4 31.7 53.4
${}^{\ddagger}$ LLaMA-13B 54.2 60.9 64.5 61.4 63.5 48.7 36.2 55.6
${}^{\ddagger}$ DeBERTa-v3-304M 57.2 64.4 69.6 72.6 62.6 57.2 42.9 60.9
${}^{\ddagger}$ flan-t5-base (248M) 56.8 58.7 72.3 64.5 68.0 54.5 41.4 59.5
flan-t5-large (783M) 62.6 58.2 76.7 67.0 72.6 58.6 48.6 63.5
flan-t5-xl (3B) 62.1 64.9 80.6 73.7 68.7 63.0 49.8 66.1
flan-t5-xxl (11B) 68.2 68.7 81.2 75.1 76.7 62.5 51.0 69.1
${}^{\ddagger}$ LLaMA-7B 63.1 68.2 75.7 69.4 69.9 68.6 47.5 66.1
Table 2: Zero-shot evaluation results, where $\dagger$ denotes IE Models and $\ddagger$ denotes NER Models. Results for ChatGPT and UniNER are from Zhou et al. (2023); InstructUIE are from Wang et al. (2023); GoLLIE are from Sainz et al. (2023); GLiNER-L are from Zaratiana et al. (2023). We bold the best results and underline the second-best results. More details about the performance including error bars are shown in Appendix E.
@@ 4.2 Instrucion Tuning
Figure 5: Prompt that are used for instruction tuning.
@@
4.3 Hierarchical Matching Algorithm
To handle the omission, addition, and substitution problems outlined in the generated predictions in our task schema, as detailed in section 3.4, we develop a Hierarchical Matching algorithm that provides a straightforward and effective solution to these challenges. Formally, given a sentence $X=\{x_{1},x_{2},\ldots,x_{n}\}$, the generated outputs can be formatted as ``$\tilde{x}_{1}(\tilde{y}_{1})\quad\tilde{x}_{2}(\tilde{y}_{2})\quad\ldots\quad\tilde{x}_{m}(\tilde{y}_{m})"$. Firstly, we utilize regular expression matching to obtain the predicted sequence $\tilde{X}=\{\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{m}\}$ and the corresponding answers $\tilde{Y}=\{\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m}\}$. Due to the inherent uncertainties in generation, $\tilde{X}$ often differs from $X$. Next, we establish a one-to-one correspondence between the words in the original sequence $X$ and the generated sequence $\tilde{X}$, then map the labels of the corresponding words $\tilde{Y}$ back to obtain the final prediction results $\hat{Y}$. A common method involves calculating the Longest Common Subsequence (LCS) between $X$ and $\tilde{X}$ to identify the correspondence between words, using the classic dynamic programming algorithm with time complexity of $O(N^{2})$. Combined with the actual NER task scenarios and our task schema, we make the following two optimizations in the matching algorithm and condition:
LCS Matching Algorithm
We optimize the complexity of the LCS algorithm using a hierarchical divide-and-conquer approach through the following steps: (1) If the sequence does not have the above problem, i.e., $\tilde{X}=X$, it is obvious that $\hat{Y}=\tilde{Y}$. The time complexity is $O(N)$, (2) For the omission case, where $\tilde{X}$ is a subsequence of $X$, the matching process can be accomplished in $O(N)$ through greedy matching. (3) In other cases, we have implemented a fast version of the LCS algorithm Hunt and Szymanski (1977) within $O(N\log N)$, based on the nature of the small number of duplicate words in $\tilde{X}$.
Back Tokenization
One notable problem in the matching process is the missing words in the vocabulary, as detailed in our case study in section C. For example, “antropología” in the original text becomes ‘antropologa’ in the model’s predictions, resulting in an inaccurate match in the matching process. To address this, we employ back tokenization, which involves tokenizing each word in the original text and then detokenizing it to match a word in the model’s vocabulary, thereby creating a more resilient matching condition.

### v2

@@ Abstract
Large Language Models (LLMs) have demonstrated impressive capabilities for generalizing in unseen tasks. In the Named Entity Recognition (NER) task, recent advancements have seen the remarkable improvement of LLMs in a broad range of entity domains via instruction tuning, by adopting entity-centric schema. In this work, we explore the potential enhancement of the existing methods by incorporating negative instances into training. Our experiments reveal that negative instances contribute to remarkable improvements by (1) introducing contextual information, and (2) clearly delineating label boundaries. Furthermore, we introduce an efficient longest common subsequence (LCS) matching algorithm, which is tailored to transform unstructured predictions into structured entities. By integrating these components, we present GNER, a Generative NER system that shows improved zero-shot performance across unseen entity domains. Our comprehensive evaluation illustrates our system’s superiority, surpassing state-of-the-art (SoTA) methods by 9 $F_{1}$ score in zero-shot evaluation. [^1] Code, datasets, and models are publicly available: https://github.com/yyDing1/GNER.
@@ 1 Introduction
To calibrate the potential enhancement of including negative instances in training, we first conduct a preliminary study. Through experiments, we show that negative instances can significantly boost the model’s performance by (1) incorporating the contextual information, and (2) enhancing the label boundary between entities and non-entities. The possible drawback of introducing the negative instances is the increase of the prediction length, leading to inaccurate predictions, reflected by the word addition, omission and substitution. To tackle the inaccuracy drawbacks, we aim to design a more accurate and efficient algorithm to convert unstructured text into structured entities.
Inspired by the above observations, we design an effective and efficient Generative NER framework named GNER. We first design a proper task schema integrating negative instances into the instruction tuning process. Additionally, we design an LCS Matching algorithm to tackle the issues in the structuring process efficiently. This innovation ensures accurate categorization and alignment of extracted entities. We also demonstrate that zero-shot performance can be enhanced with beam search through a self-correction mechanism. These strategic developments collectively advance the GNER framework, setting a new standard for accuracy and efficiency in the field of NER.
We conduct experiments on two representative generative models, Flan-T5 and LLaMA. The resulting models, GNER-T5 and GNER-LLaMA, outperform SoTA by a large margin. As stated in Fig. 1, GNER-LLaMA-7B outperforms the GoLLIE Sainz et al. (2023) trained on Code-LLaMA-7B by 6 $F_{1}$ score. Furthermore, compared to the similarly configured model UniversalNER, GNER-LLaMA-7B shows an improvement of 12.7 $F_{1}$ score, with a 2.5$\times$ boost in inference speed. We also showcase the potential of smaller models with our 780M GNER-T5-large model, which outperforms all baseline models in both zero-shot and supervised scenarios.
@@ 2 Related Work
Early works format Named Entity Recognition (NER) as a sequence labeling problem Chiu and Nichols (2016); Huang et al. (2015); Akbik et al. (2018); Qin et al. (2019); Devlin et al. (2018), utilizing the BIO-Tagging scheme. Then, different methods are proposed to address more complex scenarios, i.e., nested and discontinuous NER. These methods regard NER as question answering Li et al. (2020a); Mengge et al. (2020), span classification Fu et al. (2021); Li et al. (2020b), dependency parsing Yu et al. (2020), word-level relation classification Li et al. (2022a), and so on. In most of these approaches, negative instances have played a crucial role in the training process, either by integrating all negative instances or employing sampling methods to select part of them Li et al. (2022b). However, the performance of the above-mentioned supervised models significantly decreases in zero-shot settings Liu et al. (2021), especially when the data and domain distribution significantly diverge from those seen of training.
@@
3 Incorporating Negative Instances
In this section, we start from the entity-centric schema and explore the possible improvements by incorporating negative instances. Through experiments, we demonstrate the impact of negative instances in mitigating Unlabeled Errors (UE), Noisy Errors (NE), and Boundary Errors (BE) by (1) introducing contextual information and (2) enhancing entity boundaries.
3.1 Definition & Settings
The Named Entity Recognition (NER) task can be formally defined as a function mapping of input tokens $X=\{x_{1},x_{2},\ldots,x_{n}\}$ and a pre-defined set of entity types $L=\{l_{1},l_{2},\ldots l_{m}\}$ to entity labels $Y=\{y_{1},y_{2},\ldots,y_{n}\}$. The positive and negative instances can be formulated as follows:
@@ 3.1 Definition & Settings
where O represents non-entity text.
For the experimental setup, we choose Flan-T5-large, a model with 780M parameters, as our backbone. Additionally, we sample 10K samples from the Pile-NER Zhou et al. (2023) dataset as our training set and 200 samples for each subtask of CrossNER Liu et al. (2021) validation set to evaluate the model’s zero-shot performance. To conduct a more detailed evaluation, in addition to the $F_{1}$ score, we introduce the following three metrics to assess model performance:
Unlabeled Error (UE)
The model fails to recognize the entity and labels it as “O”.
Noisy Error (NE)
The model mistakenly label an entity with another incorrect entity tag.
Boundary Error (BE)
The model correctly predicts the entity type but fails to identify its full extent, either capturing only a portion of the entity or resulting in overlaps.
An example of the constructed prompt Token inputs ($X$): John explored Tokyo , sampling its famed sushi , and flew back to New York . Entity type ($L$): [Person, Location]   Training prompt w/o context (entity-centric): [John](Person) [Tokyo](Location) [New York](Location) w/ context length 1: [John](Person) explored [Tokyo](Location) , …… to [New York](Location) . w/ full context: [John](Person) explored [Tokyo](Location) , sampling its famed sushi , and flew back to [New York](Location). w/ full context and label boundary (BIO): John(B-Person) explored(O) Tokyo(B-Location) ,(O) sampling(O) its(O) famed(O) sushi(O) ,(O) and(O) flew(O) back(O) to(O) New(B-Location) York(I-Location) .(O) w/ full context and label boundary (BIOES): John(S-Person) explored(O) Tokyo(S-Location) ,(O) sampling(O) its(O) famed(O) sushi(O) ,(O) and(O) flew(O) back(O) to(O) New(B-Location) York(I-Location) .(O)
Figure 3: Constructed prompts used for training.
UE NE BE $F_{1}$
w/o context (entity-centric) 7.8 16.7 3.8 59.0
w/ context length 1 7.6 15.7 3.8 60.3
w/ full context 7.7 14.9 3.7 61.0
w/ full context & BIO 7.5 14.4 3.3 61.8
w/ full context & BIOES 7.6 14.7 3.5 61.2
Table 1: Unlabeled Error (UE), Noisy Error (NE) and Boundary Error (BE) in our preliminary study.
@@ 3.2 Learning with Entity Context
We integrate the contextual information before and after an entity into our training process to explicitly enable the model to recognize entities based on their surrounding context. Specifically, we introduce negative instances that are closest to the entity, extending up to a length $L$, until encountering the boundary of the sentence, as part of our training instances. An example of our constructed training prompt is shown in Fig. 3, and the corresponding results are summarized in Table 1 and Fig. 4.
The context surrounding entities plays a significant role in determining their categories, with a notable improvement observed when increasing the contextual length from 0 to 1. As the contextual length increases, performance progressively improves, showing that the model benefits from the context, and the ratio of noisy error (NE) significantly decreases, largely contributing to the improvement in the final $F_{1}$ score. Qualitatively, we further analyze through case studies why the closest negative instances contribute to improvement. We discover that the model tends to learn more from the context, such as “flew to” prompting the model to focus more on the following entity “New York” instead of merely memorizing “New York” as a location. We also experiment by placing terms with multiple meanings, such as Jordan, Amazon, and Mercury, after “flew to” and observe that the model consistently identified them correctly.
@@ 3.3 Entity Boundary of Generative Model
The above analysis has demonstrated the effectiveness of entity contexts, where labels are applied exclusively to entity parts. We then adopt BIO-tagging Huang et al. (2015) to enrich all label information, with the training prompt as shown in Fig. 3. The results in Table 1 indicate that the introduction of BIO tagging effectively strengthens the boundaries around entities, leading to a significant reduction in boundary errors. As shown in Fig. 4, compared to training with contexts alone, incorporating the BIO-tagging strategies results in consistent improvements across various contextual lengths. In addition to BIO tagging, we also try the BIOES tagging scheme, based on the intuition that the BIOES tagging method provides a stronger delineation of entity boundaries. However, we find that the performance of BIOES tagging is not as good as BIO tagging. The results are listed in Table 1. Upon further analysis, we discover that the BIOES tagging seems harder to learn under the auto-regressive generation manner, where each token is predicted sequentially. For instance, both “B-” and “S-” can serve as the beginning of an entity, but only “B-” can be followed by “I-”, which may confuse the model for subsequent words.
3.4 The role of negative instances
The improvement in model performance ($F_{1}$ score) can be explained as follows:
Precision and Recall
To put it more directly, the improvements over entity-centric approaches are primarily reflected in (1) Precision: The context surrounding an entity often leads to a more accurate determination of its type, and (2) Recall: The model is guided to make judgments on every token in a sentence (including those in non-entity texts), which helps recall more entities. We observe improvements in both recall and precision, which in turn lead to an increase in the $F_{1}$ score.
Less Unlabeled, Noisy and Boundary Error
As indicated in Table 1, context helps mitigate Unlabeled and Noisy Errors, while BIO tagging strengthens entity boundaries, thereby reducing Boundary Errors. These reductions in errors directly contribute to improved performance.
@@ 4 Method
In this section, we present our GNER framework. We start by describing our task schema, which integrates negative instances into the training process for better usage of contextual information (section 3.2) and sensitivity to the entity boundaries (section 3.3), followed by the correlated tuning strategies. We also propose an effective longest common subsequence (LCS) matching algorithm, to convert the model’s unstructured text outputs into structured data efficiently, thereby enhancing the accuracy of our system.
@@ 4.2 Instrucion Tuning
Figure 5: Prompt used for instruction tuning.
@@
4.3 LCS Matching Algorithm
The prediction length of the model increases with the integration of entity contexts and BIO tagging strategies. A longer generation sequence might bring challenges to the popular generative LLMs. The model’s output may include omissions, additions, and substitutions of words. We launch a detailed case study and find the potential causes of these issues: (1) noise in the original text, (2) missing words in the vocabulary, and (3) accumulative exposure bias. The representative examples, issue proportion, along with detailed analysis, are documented in Appendix C.
To handle these problems, we develop a LCS Matching algorithm that provides a straightforward and effective solution to these challenges. Formally, given a sentence $X=\{x_{1},x_{2},\ldots,x_{n}\}$, the generated outputs can be formatted as ``$\tilde{x}_{1}(\tilde{y}_{1})\quad\tilde{x}_{2}(\tilde{y}_{2})\quad\ldots\quad\tilde{x}_{m}(\tilde{y}_{m})"$. Firstly, we utilize regular expression matching to obtain the predicted sequence $\tilde{X}=\{\tilde{x}_{1},\tilde{x}_{2},\ldots,\tilde{x}_{m}\}$ and the corresponding answers $\tilde{Y}=\{\tilde{y}_{1},\tilde{y}_{2},\ldots,\tilde{y}_{m}\}$. Due to the inherent uncertainties in generation, $\tilde{X}$ often differs from $X$. Next, we establish a one-to-one correspondence between the words in the original sequence $X$ and the generated sequence $\tilde{X}$, then map the labels of the corresponding words $\tilde{Y}$ back to obtain the final prediction results $\hat{Y}$. A common method involves calculating the Longest Common Subsequence (LCS) between $X$ and $\tilde{X}$ to identify the correspondence between words, using the classic dynamic programming algorithm with time complexity of $O(N^{2})$. Combined with the actual NER task scenarios and our task schema, we make the optimizations in the matching algorithm and condition. The resulting algorithm can handle these issues effectively with a time complexity of $O(N\log N)$. The optimization in matching conditions (i.e., Back Tokenization) further enhances the robustness across various models in our system. More details concerning the optimization can be seen in Appendix D.